A. Prove that Km equals the substrate concentration at ½ Vmax .

(Show solution)

B. Why do structural analogs if the transition-state intermediate of an enzyme inhibit the enzyme competitively and with low Ki values (binds tightly)?
Answer #1

MICHAELIS-MENTEN KINETICS AND THE STEADY STATE Review Enzymes make reactions Golfastes. E+s 1ES 23 E+ P Rate = k, [E] [s) Rat

le, +62 km (€)6) = (es) em + (ES) (5) ces) (ling + (s)) - CE) (5) (es) - Cer(s) +(s) Vocale = 60,.(65) (Goles): 62.[6), C5) k